OpenStudy (anonymous):

hich of the following is the solution to the equation 36^3c = square root of 6^(6c + 8)

jhonyy9 (jhonyy9):

can you rewriting every two sides based on 6 on the respectly exponents ?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

\[36^{3c}=\sqrt{6}^{6c-8}\]

jhonyy9 (jhonyy9):

yes but based on 6 on the needed exponts like 36=6^2 for example

OpenStudy (anonymous):

Do you know what the answer is cause your're just confusing me more

jhonyy9 (jhonyy9):

so 36^3c =(6^2)^3c =6^6c ok ? can you rewriting the right side in this form ?

jhonyy9 (jhonyy9):

came en i like help you understanding how you can solve it easy

jhonyy9 (jhonyy9):

how you write sqrt6 using exponent form ?

jhonyy9 (jhonyy9):

you dont like cooperating ?

OpenStudy (anonymous):

im doing a flvs seg 2 exam so im trying to that as well

jhonyy9 (jhonyy9):

6^6c =6^((6c+8)/2)

jhonyy9 (jhonyy9):

6^6c =6^(3c+4) than now make equale these exponents 6c=3c+4 3c=4 c=4/3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

thank you very much

jhonyy9 (jhonyy9):

yw good luck bye